# Explaining Explainability: Understanding Concept Activation Vectors

### v1

@@ Abstract
Recent interpretability methods propose using concept-based explanations to translate the internal representations of deep learning models into a language that humans are familiar with: concepts. This requires understanding which concepts are present in the representation space of a neural network. One popular method for finding concepts is Concept Activation Vectors (CAVs), which are learnt using a probe dataset of concept exemplars. In this work, we investigate three properties of CAVs. CAVs may be: (1) inconsistent between layers, (2) entangled with different concepts, and (3) spatially dependent. Each property provides both challenges and opportunities in interpreting models. We introduce tools designed to detect the presence of these properties, provide insight into how they affect the derived explanations, and provide recommendations to minimise their impact. Understanding these properties can be used to our advantage. For example, we introduce spatially dependent CAVs to test if a model is translation invariant with respect to a specific concept and class. Our experiments are performed on ImageNet and a new synthetic dataset, Elements. Elements is designed to capture a known ground truth relationship between concepts and classes. We release this dataset to facilitate further research in understanding and evaluating interpretability methods.
Keywords:
Interpretability Explainable AI Concepts
@@ 1 Introduction
Deep learning models have become ubiquitous, achieving performance reaching or surpassing human experts across a variety of tasks. However, currently, the inherent complexity of these models obfuscates our ability to explain their decision-making process. As they are applied in a growing number of real-world domains, there is an increasing need to understand how they work. This transparency allows for easier debugging and better understanding of model limitations.
Model explanations can take many forms, such as input features, prototypes or concepts. Recent work has shown that explainability methods that focus on low-level features can incur problems. For example, saliency methods can suffer from confirmation bias and lack faithfulness [2]. Even when faithful, they only show ‘where’ the model focused in the image, and not ‘what’ it focused on [1, 9].
To address these problems, concept-based methods provide explanations using high-level terms that humans are familiar with. A popular method is concept activation vectors (CAVs): a linear representation of a concept found in the activation space of a specific layer using a probe dataset of concept examples [23]. However, concept-based methods also face challenges, such as their sensitivity to the specific probe dataset [34, 38].
@@
We provide tools to analyse each property and show that they can affect testing with CAVs (TCAV) (§6.1, §6.2 and §6.3). To minimise the impact these effects can have, we recommend: creating CAVs for multiple layers, verifying expected dependencies between related concepts, and visualising spatial dependence (§7). These properties do not imply that CAVs should not be used. On the contrary, we may be able to use these properties to better understand model behaviour. For example, we introduce a modified version of CAVs that are spatially dependent and can be used to identify translation invariance in convolutional neural networks (CNNs).
@@
Figure 1: Concept Activation Vectors can be: inconsistent, i.e., we cannot find two concept vectors in different layers that have the same additive effect (left), entangled (middle) and spatially dependent (right). The top panel illustrates each of these different properties. The bottom panels show our recommendations on how to minimise the impact these effects can have: creating CAVs for multiple layers (left), verifying expected dependencies between related concepts (middle), and visualising spatial dependence (right).
@@ 2 Background: Concept Activation Vectors
A CAV [23] is a vector representation of a concept found in the activation space of a layer of a NN. Consider a NN which can be decomposed into two functions: $g_{l}({\bm{x}})={\bm{a}}_{l}\in\mathbb{R}^{m}$ which maps the input ${\bm{x}}\in\mathbb{R}^{n}$ to a vector ${\bm{a}}_{l}$ in the activation space of layer $l$, and $h_{l}({\bm{a}}_{l})$ which maps ${\bm{a}}_{l}$ to the output. To create a CAV for a concept $c$ we need a probe dataset $\mathbb{D}_{c}$ consisting of positive samples $\mathbb{X}_{c}^{+}$ (concept examples), and negative samples $\mathbb{X}_{c}^{-}$ (random in-distribution images). For the sets $\mathbb{X}_{c}^{-}$ and $\mathbb{X}_{c}^{+}$, we create a corresponding set of activations in layer $l$:
@@
To analyse a model’s sensitivity to ${\bm{v}}_{c,l}$, Kim et al. [23] introduce testing with CAVs (TCAV), which determines the model’s conceptual sensitivity across an entire class. Let $\mathbb{X}_{k}$ be a set of inputs belonging to class $k$. The TCAV score is defined as
@@
where the directional derivative of the concept, $S_{c,k,l}$, defined as
@@
where $\nabla h_{l,k}$ is the partial derivative of the NN output for class $k$ to the activation. The TCAV score measures the fraction of class $k$ inputs whose activation at layer $l$ is positively influenced by concept $c$. A statistical test comparing the scores of CAVs to random vectors is used to determine the concept’s significance (see Appendix 11.1).
@@ 3 CAV Hypotheses
3.1 Consistency
In general, we want to understand model behaviour. However, CAVs explain whether a model is sensitive to a concept in a specific layer. In practice, analysing all layers may be computationally infeasible, and it is unclear which layers to choose. Therefore, our first hypothesis explores the relationship between CAVs found in different layers. Recall that the TCAV scores depend on the directional derivative: how the model output changes for an infinitesimal change of the activations in the direction of a CAV. By perturbing the activations in the direction of a CAV, we explore whether two concept vectors found in different layers can have the same affect on the model output. We refer to this property as consistency (see Figure 1 for a schematic overview).
Definition 1 (consistent)
Assume we have a function $f(\cdot)$ that maps the activations from layer $l_{1}$ into activations in layer $l_{2}$, where $l_{1}<l_{2}$. Concept vectors, ${\bm{v}}_{c,l_{1}}$ and ${\bm{v}}_{c,l_{2}}$ are consistent iff for every input ${\bm{x}}$ and corresponding activations ${\bm{a}}_{l_{1}}$ and ${\bm{a}}_{l_{2}}$, $f({\bm{a}}_{l_{1}}+{\bm{v}}_{c,l_{1}})={\bm{a}}_{l_{2}}+{\bm{v}}_{c,l_{2}}$.
If two CAVs are consistent then they have the same downstream affect on the model when activations are perturbed in their direction, i.e., even though they are in different layers, they have an equivalent effect on the model output and therefore the model assigns them the same meaning. Our first hypothesis is:
@@ 3.1 Consistency
In §6.1 we formally explore this hypothesis, and perform empirical evaluations on the Elements and ImageNet [12] datasets. We show theoretically the conditions ${\bm{v}}_{c,l_{2}}$ and ${\bm{a}}_{l_{1}}$ must meet for vectors ${\bm{v}}_{c,l_{1}}$ and ${\bm{v}}_{c,l_{2}}$ to be consistent when $f$ is either a rectified linear unit (ReLU) or sigmoid function.
@@ A CAV ${\bm{v}}_{c_{1},l}$ for concept $c_{1}$ is entangled with concept $c_{2}$ iff
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{a}}_{c_{2},l}^{+}}\cdot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\bm{v}}_{c_{1},l}}>{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{a}}_{c_{2},l}^{-}}\cdot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\bm{v}}_{c_{1},l}}$ $\forall{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{a}}_{c_{2},l}^{+}}\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{A}_{c_{2},l}^{+}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{a}}_{c_{2},l}^{-}}\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{A}_{c_{2},l}^{-}}$ (6)
@@
In §6.2 we provide a visualisation tool to explore CAV entanglement and discuss how this can affect TCAV.
@@ 3.3 Spatial Dependence
Here, we explore the influence of spatial dependence on concepts. Let $\mathbb{D}_{c,\mu_{1}}$ and $\mathbb{D}_{c,\mu_{2}}$ denote two datasets containing the same concept but in different locations $\mu_{1}\neq\mu_{2}$. For example, $\mathbb{D}_{c,\mu_{1}}$ may contain exemplars of striped on the left of the image, and $\mathbb{D}_{c,\mu_{2}}$ exemplars of the striped on the right of the image – an example is shown in Fig. 2. As before, we construct latent representations $\mathbb{A}_{c,l,\mu_{1}}$ and $\mathbb{A}_{c,l,\mu_{2}}$ for datasets $\mathbb{D}_{c,\mu_{1}}$ and $\mathbb{D}_{c,\mu_{2}}$, respectively. Let ${\bm{v}}_{c,l}$ be the concept vector found using probe dataset $\mathbb{D}_{c,\mu_{1}}$.
@@
Let ${\bm{a}}_{l,i}$ be the activations corresponding to input ${\bm{x}}_{i}$ in layer $l$, and let $\mu_{c,i}$ be the location of concept $c$ in ${\bm{x}}_{i}$. A layer has a spatially dependent representation of a concept iff
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\bm{a}}_{c,l,\mu_{1}}^{+}}\cdot{\bm{v}}_{c,l}>{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{a}}_{c,l,\mu_{2}}^{+}}\cdot{\bm{v}}_{c,l}$ $\forall{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\bm{a}}_{c,l,\mu_{1}}^{+}}\in{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{A}_{c,l,\mu_{1}}^{+}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bm{a}}_{c,l,\mu_{2}}^{+}}\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{A}_{c,l,\mu_{2}}^{+}}.$ (8)
@@ 5 Related Work
Chen et al. [7] discuss how concept vectors can be correlated, making it challenging to create a vector that solely represents one concept. While their work focuses on de-correlating concepts during training, we focus on analysing the impact of correlated concepts after training and show how they can lead to misleading explanations (§6.2). Fong and Vedaldi [15] use cosine similarity to demonstrate that the similarity between concepts varies based on the vector creation method. In our work, we also use cosine similarity to compare concept vectors. The distinction lies in our focus on CAVs and the insights they provide into the dataset and model.
@@
Biscione and Bowers [6] describe how CNNs are not inherently translation invariant but can learn to be (under certain conditions on the dataset). This finding challenges the common assumption that CNNs possess inherent translation invariance. Through spatially dependent CAVs, we demonstrate translation invariance with respect to a specific concept and class, rather than in general, providing more detailed information about a model.
@@
Most concept-based interpretability methods represent concepts as vectors in the activation space of a trained neural network (NN) [23, 15, 50, 17, 48, 35, 14]. However, some concept-based methods use different representations: individual neurons [5], regions of activation space [11] or non-linear concepts [4, 26]. Our work focuses on the properties of concept vectors.
@@
To give insight into when the various properties may be relevant, we performed a review of computer vision papers which use CAVs in (1) the high-stakes applications of medical imaging (including skin cancer, skin lesions, breast cancer, and histology [42, 16, 33]), and (2) computer vision research on models trained with well-known datasets [25, 27, 40, 49, 36, 12]. A summary table can be found in Appendix 16. We found that the following papers could have benefited from evaluating: consistency [42, 34, 16, 45, 18, 28], entanglement [42, 34, 16, 45, 18, 19, 29, 28, 33], and spatial dependence [42, 34, 16, 45, 18, 29, 28, 33]. We provide a detailed example, using the application of skin cancer diagnosis [42], in § 7 and Appendix 17.
@@
While several datasets have been introduced for evaluating interpretability methods, they differ from ours in a few key ways. There are three aspects we care about:
@@
Existing datasets only allow insight into (1), whereas our dataset allows us to analyse (2) and (3) as well. The Benchmarking Interpretability Method (BIM) [43] inserts objects into scene images. While it benefits from utilizing real images and complex concepts (dog or bedroom), it also presents challenges. One drawback is that relying on real images makes it challenging to establish the ground truth relationship between concepts and class predictions or to know the similarities between concepts. As such, it does not give us insight into (2) or (3). The synthetic dataset in Yeh et al. [44] is the most similar to our dataset but was designed for concept discovery, featuring images where each object corresponds to a single concept (shape). In our dataset, each object contains multiple concepts, allowing us to create associations between them. We focus on explanation faithfulness by ensuring that the concepts must be used correctly by the model to achieve a high accuracy. So, for an accurate model, we have a ground truth understanding of how each concept is used. An extended literature review can be found in Appendix 16.

### v2

# Explaining Explainability: Recommendations for Effective Use of Concept Activation Vectors

@@ Abstract
Concept-based explanations translate the internal representations of deep learning models into a language that humans are familiar with: concepts. One popular method for finding concepts is Concept Activation Vectors (CAVs), which are learnt using a probe dataset of concept exemplars. In this work, we investigate three properties of CAVs: (1) inconsistency across layers, (2) entanglement with other concepts, and (3) spatial dependency. Each property provides both challenges and opportunities in interpreting models. We introduce tools designed to detect the presence of these properties, provide insight into how each property can lead to misleading explanations, and provide recommendations to mitigate their impact. To demonstrate practical applications, we apply our recommendations to a melanoma classification task, showing how entanglement can lead to uninterpretable results and that the choice of negative probe set can have a substantial impact on the meaning of a CAV. Further, we show that understanding these properties can be used to our advantage. For example, we introduce spatially dependent CAVs to test if a model is translation invariant with respect to a specific concept and class. Our experiments are performed on natural images (ImageNet), skin lesions (ISIC $2019$), and a new synthetic dataset, Elements. Elements is designed to capture a known ground truth relationship between concepts and classes. We release this dataset to facilitate further research in understanding and evaluating interpretability methods.
@@ 1 Introduction
Deep learning models have become ubiquitous, achieving performance reaching or surpassing human experts across a variety of tasks. However, currently, the inherent complexity of these models obfuscates our ability to explain their decision-making process. As they are applied in a growing number of real-world domains, there is an increasing need to understand how they work (Doshi-Velez & Kim, 2017; Reyes et al., 2020). This transparency allows for easier debugging and better understanding of model limitations.
Model explanations can take many forms, such as input features, prototypes or concepts. Recent work has shown that explainability methods that focus on low-level features can incur problems. For example, saliency methods, which determine the sensitivity of a deep learning model to individual pixels, can suffer from confirmation bias and lack model faithfulness (Adebayo et al., 2018). Even when faithful, saliency maps only show ‘where’ the model focused in the image, and not ‘what’ it focused on (Achtibat et al., 2022; Colin et al., 2022).
To address these problems, concept-based methods provide explanations using high-level terms that humans are familiar with. A popular method is concept activation vectors (CAVs): a linear representation of a concept found in the activation space of a specific layer using a probe dataset of concept examples (Kim et al., 2018). However, concept-based methods also face challenges, such as their sensitivity to the specific probe dataset (Ramaswamy et al., 2022a; Soni et al., 2020).
@@
We provide tools to analyse each property and show that they can affect testing with CAVs (TCAV) (§6.1, §6.2 and §6.3) and lead to misleading explanations. To minimise the impact these effects can have, we recommend: creating CAVs for multiple layers, verifying expected dependencies between related concepts, and visualising spatial dependence (§7). These properties do not imply that CAVs should not be used. On the contrary, we may be able to use these properties to better understand model behaviour. For example, we introduce a modified version of CAVs that are spatially dependent and can be used to identify translation invariance in convolutional neural networks (CNNs). Additionally, we do not claim that these properties are necessarily unexpected. Instead, the aim of this paper is to formally define each one, confirm when it can occur, and then clearly examine the consequences of these properties on the interpretation of TCAV scores. In each case, even when expected, these properties can cause misleading explanations.
To provide a concrete example of how the properties can affect experiments, we examine the use-case of Yan et al. (2023) which uses CAVs in the context of skin cancer diagnosis (§ 7). We demonstrate how to use our recommendations to sanity-check TCAV results and, in this specific case, we show that entangled concepts lead to uninterpretable explanations. Our results also indicate that the choice of negative probe dataset can have a substantial impact to the meaning of a CAV. This use-case, combined with our recommendations, can be used by practitioners as a guide for how to effectively use CAVs in practice.
@@
Figure 1: Concept Activation Vectors can be: inconsistent across layers, i.e., we cannot find two concept vectors in different layers that have the same additive effect (left), entangled (middle) and spatially dependent (right). The top panel illustrates each of these different properties. The bottom panels show our recommendations on how to mitigate the impact these effects can have: creating CAVs for multiple layers (left), verifying expected dependencies between related concepts (middle), and visualising spatial dependence (right).
@@ 2 Background: Concept Activation Vectors
A CAV (Kim et al., 2018) is a vector representation of a concept found in the activation space of a layer of a neural network (NN). Consider a NN which can be decomposed into two functions: $g_{l}({\bm{x}})={\bm{a}}_{l}\in\mathbb{R}^{m}$ which maps the input ${\bm{x}}\in\mathbb{R}^{n}$ to a vector ${\bm{a}}_{l}$ in the activation space of layer $l$, and $h_{l}({\bm{a}}_{l})$ which maps ${\bm{a}}_{l}$ to the output. To create a CAV for a concept $c$ we need a probe dataset $\mathbb{D}_{c}$ consisting of positive samples $\mathbb{X}_{c}^{+}$ (concept examples), and negative samples $\mathbb{X}_{c}^{-}$ (randomly sampled in-distribution images). For the sets $\mathbb{X}_{c}^{-}$ and $\mathbb{X}_{c}^{+}$, we create a corresponding set of activations in layer $l$:
@@
To analyse a model’s sensitivity to ${\bm{v}}_{c,l}$, Kim et al. (2018) introduce testing with CAVs (TCAV), which determines the model’s conceptual sensitivity across an entire class. Let $\mathbb{X}_{k}$ be a set of inputs belonging to class $k$. The TCAV score is defined as
@@
where the directional derivative of the concept, $S_{c,k,l}$, is defined as
@@
where $\nabla h_{l,k}$ is the partial derivative of the NN output for class $k$ with respect to the activation. The TCAV score measures the fraction of class $k$ inputs whose activation at layer $l$ is positively influenced by concept $c$. A statistical test comparing the scores of CAVs to random vectors is used to determine the concept’s significance (see Appendix 10.1).
@@ 3 CAV Hypotheses
3.1 Layer Consistency
In general, we want to understand model behaviour. However, CAVs explain whether a model is sensitive to a concept in a specific layer. In practice, analysing all layers may be computationally infeasible, and it is unclear which layers to choose. Therefore, our first hypothesis explores the relationship between CAVs found in different layers. Recall that the TCAV scores depend on the directional derivative: how the model output changes for an infinitesimal change of the activations in the direction of a CAV. By perturbing the activations in the direction of a CAV, we explore whether two concept vectors found in different layers can have the same affect on the model output. We refer to this property as layer consistency (see Figure 1 for a schematic overview).
Definition 1 (layer consistency)
Assume we have a function $f(\cdot)$ that maps the activations from layer $l_{1}$ into activations in layer $l_{2}$, where $l_{1}<l_{2}$. Concept vectors, ${\bm{v}}_{c,l_{1}}$ and ${\bm{v}}_{c,l_{2}}$ are consistent across layers iff for every input ${\bm{x}}$ and corresponding activations ${\bm{a}}_{l_{1}}$ and ${\bm{a}}_{l_{2}}$, $f({\bm{a}}_{l_{1}}+{\bm{v}}_{c,l_{1}})={\bm{a}}_{l_{2}}+{\bm{v}}_{c,l_{2}}$.
If two CAVs are consistent across layers then they have the same downstream affect on the model when activations are perturbed in their direction, i.e., even though they are in different layers, they have an equivalent effect on the model output and therefore the model assigns them the same meaning.  [^2] For simplicity we write the concept vectors without a scaling term, but in all experiments we scale by some small $\gamma$ and the mean norm of the activations in that layer – see Appendix 12.1 for details. Our first hypothesis is:
@@ 3.1 Layer Consistency
In §6.1 we formally explore this hypothesis, and perform empirical evaluations on the Elements and ImageNet (Deng et al., 2009) datasets. We show theoretically the conditions $f$ must meet for layer consistent vectors ${\bm{v}}_{c,l_{1}}$ and ${\bm{v}}_{c,l_{2}}$ to exist.
@@ A CAV ${\bm{v}}_{c_{1},l}$ for concept $c_{1}$ is entangled with concept $c_{2}$ iff
${\color[rgb]{0,0,1}{\bm{a}}_{c_{2},l}^{+}}\cdot{\color[rgb]{1,0,0}{\bm{v}}_{c_{1},l}}>{\color[rgb]{0,0,1}{\bm{a}}_{c_{2},l}^{-}}\cdot{\color[rgb]{1,0,0}{\bm{v}}_{c_{1},l}}$ $\forall{\color[rgb]{0,0,1}{\bm{a}}_{c_{2},l}^{+}}\in{\color[rgb]{0,0,1}\mathbb{A}_{c_{2},l}^{+}},{\color[rgb]{0,0,1}{\bm{a}}_{c_{2},l}^{-}}\in{\color[rgb]{0,0,1}\mathbb{A}_{c_{2},l}^{-}}$ (6)
@@
In §6.2 we demonstrate that visualising similarity matrices can be used to explore CAV entanglement and discuss how entanglement can affect TCAV.
@@ 3.3 Spatial Dependence
Here, we explore the influence of spatial dependence on concepts. Let $\mathbb{D}_{c,\mu_{1}}$ and $\mathbb{D}_{c,\mu_{2}}$ denote two datasets containing the same concept but in different locations $\mu_{1}\neq\mu_{2}$. By location we mean the location of the concept relative to the frame of the image. The exact form of this will depend on the specific dataset and concept in question. For example, in the Elements dataset we use binary labels such as left/right or top/bottom but more complex representations could be used (e.g. a segmentation map of which pixels contain a concept). For example, $\mathbb{D}_{c,\mu_{1}}$ may contain exemplars of striped on the left of the image, and $\mathbb{D}_{c,\mu_{2}}$ exemplars of the striped on the right of the image – an example is shown in fig. 2. As before, we construct latent representations $\mathbb{A}_{c,l,\mu_{1}}$ and $\mathbb{A}_{c,l,\mu_{2}}$ for datasets $\mathbb{D}_{c,\mu_{1}}$ and $\mathbb{D}_{c,\mu_{2}}$, respectively. Let ${\bm{v}}_{c,l}$ be the concept vector found using probe dataset $\mathbb{D}_{c,\mu_{1}}$.
@@
Let ${\bm{a}}_{l,i}$ be the activations corresponding to input ${\bm{x}}_{i}$ in layer $l$, and let $\mu_{c,i}$ be the location of concept $c$ in ${\bm{x}}_{i}$. A layer has a spatially dependent representation of a concept iff there exists some function $\phi$ which maps ${\bm{a}}_{l,i}$ to $\mu_{c,i}$ for all inputs ${\bm{x}}_{i}$:
@@
${\color[rgb]{1,0,0}{\bm{a}}_{c,l,\mu_{1}}^{+}}\cdot{\bm{v}}_{c,l}>{\color[rgb]{0,0,1}{\bm{a}}_{c,l,\mu_{2}}^{+}}\cdot{\bm{v}}_{c,l}$ $\forall{\color[rgb]{1,0,0}{\bm{a}}_{c,l,\mu_{1}}^{+}}\in{\color[rgb]{1,0,0}\mathbb{A}_{c,l,\mu_{1}}^{+}},{\color[rgb]{0,0,1}{\bm{a}}_{c,l,\mu_{2}}^{+}}\in{\color[rgb]{0,0,1}\mathbb{A}_{c,l,\mu_{2}}^{+}}.$ (8)
@@ 5 Related Work
Chen et al. (2020) discuss how concept vectors can be correlated, making it challenging to create a vector that solely represents one concept. While their work focuses on de-correlating concepts during training, we focus on analysing the impact of correlated concepts after training and show how they can lead to misleading explanations (§6.2). Fong & Vedaldi (2018) use cosine similarity to demonstrate that the similarity between concepts varies based on the vector creation method. In our work, we also use cosine similarity to compare concept vectors. The distinction lies in our focus on CAVs and the insights they provide into the dataset and model. There have been several works analysing correlated concepts in interpretable-by-design networks (Heidemann et al., 2023; Zarlenga et al., 2023). Our work complements these works by studying standard neural network architectures and the post-hoc explanations of TCAV. Raman et al. (2024) examine the effect of inter-concept relationships in CAVs and provide several useful metrics for measuring how well these relationships are represented, but they do not explore its effect on TCAV scores.
@@
Biscione & Bowers (2021) describe how CNNs are not inherently translation invariant but can learn to be (under certain conditions on the dataset). This finding challenges the common assumption that CNNs possess inherent translation invariance. Through spatially dependent CAVs, we demonstrate translation invariance with respect to a specific concept and class, rather than in general, providing more detailed information about a model. Raman et al. (2023) examine the locality of concept bottleneck models (CBMs) and find that, in some cases, CBMs make concept predictions using information far from the object of interest. In this work we examine post-hoc concept-based explanations, rather than predictions of an interpretable-by-design model. Additionally, our definition of spatial dependence is more related to whether we can learn CAVs which mean stripes on the right of the image, rather than whether a CAV is using information near/far from where the concept is located in the image.
@@
Most concept-based interpretability methods represent concepts as vectors in the activation space of a trained neural network (Kim et al., 2018; Fong & Vedaldi, 2018; Zhou et al., 2018; Ghorbani et al., 2019; Zhang et al., 2020; Ramaswamy et al., 2022b; Fel et al., 2023). However, some concept-based methods use different representations: individual neurons (Bau et al., 2017), regions of activation space (Crabbé & van der Schaar, 2022) or non-linear concepts (Bai et al., 2022; Li et al., 2023). Our work focuses on the properties of concept vectors.
@@
To give insight into when the various properties may be relevant, we performed a review of computer vision papers which use CAVs in (1) the high-stakes applications of medical imaging (including skin cancer, skin lesions, breast cancer, and histology (Yan et al., 2023; Fürböck et al., 2022; Pfau et al., 2020)), and (2) computer vision research on models trained with well-known datasets (Krizhevsky, 2009; Lin et al., 2014; Wah et al., 2011; Zhou et al., 2017; Sagawa et al., 2020; Deng et al., 2009). A summary table can be found in Appendix 15. We found that the following papers could have benefited from evaluating: consistency (Yan et al., 2023; Ramaswamy et al., 2022a; Fürböck et al., 2022; Yuksekgonul et al., 2023; Ghosh et al., 2023; Lucieri et al., 2020), entanglement (Yan et al., 2023; Ramaswamy et al., 2022a; Fürböck et al., 2022; Yuksekgonul et al., 2023; Ghosh et al., 2023; Graziani et al., 2020; McGrath et al., 2022; Lucieri et al., 2020; Pfau et al., 2020), and spatial dependence (Yan et al., 2023; Ramaswamy et al., 2022a; Fürböck et al., 2022; Yuksekgonul et al., 2023; Ghosh et al., 2023; McGrath et al., 2022; Lucieri et al., 2020; Pfau et al., 2020). We provide a detailed example, using the application of skin cancer diagnosis (Yan et al., 2023), in § 7.
@@
While several datasets have been introduced for evaluating interpretability methods, they differ from ours in a few key ways. There are three questions we need our dataset to help answer:
@@
Existing datasets either do not allow insight into all three, or they have other practical reasons for being unsuitable. The Benchmarking Interpretability Method (BIM) (Yang & Kim, 2019) inserts objects into scene images. While it benefits from utilizing real images and complex concepts (dog or bedroom), it also presents challenges. One drawback is that relying on real images makes it challenging to establish the ground truth relationship between concepts and class predictions or to know the similarities between concepts. As such, it does not give us insight into (2) or (3). The CLEVR dataset (Johnson et al., 2016) could give insight into all $3$, but because it renders $3$D shapes it is too slow for our purposes. Elements generates images in $0.004$\mathrm{s}$$ compared to CLEVR’s $4$\mathrm{s}$$. This translates to a significant time saving when more data is required – $4$\mathrm{s}$$ for $1000$ images with Elements versus $1$\mathrm{h}$$ for CLEVR. Analyzing CAV properties is our core focus. Elements, with its speed and flexibility, allows us to create the many different dataset versions required for experiments. The Navon and Trifeature datasets, used by Hermann & Lampinen (2020) to study feature representations, could also give insight into the three questions, with associated concepts of shape, color and texture relating to each image. However, there is only one large object in each image so our experiments on spatial dependence would not have been possible. The synthetic dataset in Yeh et al. (2020) is similar to our dataset but it was designed for concept discovery, featuring images where each object corresponds to a single concept (shape). In our dataset, each object contains multiple concepts, allowing us to create associations between them. We focus on explanation faithfulness by ensuring that the concepts must be used correctly by the model to achieve a high accuracy. So, for an accurate model, we have a ground truth understanding of how each concept is used. dSprites Matthey et al. (2017) and 3D Shapes Burgess & Kim (2018) are probably the most similar datasets to ours but Elements is a far larger dataset. For example, even for the simple version of the Elements dataset, there are approximately $10^{10}$ different images after just a single object has been placed in the image. This is considerably larger than the $737,280$ and $480,000$ total images in the dSprites and 3D Shapes datasets, respectively. The additional complexity of Elements means an NN needs to approximate a more complex function and having multiple objects per image ensures the model has to learn object/location-based representations of the concepts, rather than image-based, e.g., to answer whether a striped triangle is present it is not sufficient to simply determine if stripes occur anywhere in the image. An extended literature review can be found in Appendix 15.